# OpenRT: An Open-Source Red Teaming Framework for Multimodal LLMs

### v1

@@ 3.2.3 Attack
Table 2 classifies attacks by strategy paradigm, including gradient optimization, genetic algorithms, LLM-driven refinement, linguistic encoding, contextual deception, multimodal exploitation, and multi-agent coordination. White-box attacks include gradient-based methods like GCG zou2023universal and Visual Jailbreak qi2024visual. Black-box attacks span multiple categories: optimization and fuzzing approaches (AutoDAN liu2023autodan, GPTFuzzer yu2023gptfuzzer), LLM-driven refinement (PAIR chao2023jailbreaking, DrAttack li2024drattack), linguistic and encoding manipulations (CipherChat yuan2023gpt, CodeAttack ren2024codeattack, FlipAttack liu2024flipattack), contextual deception strategies (Crescendo russinovich2025great, RedQueen jiang2024red), multimodal-specific attacks (FigStep gong2025figstep, IDEATOR wang2025ideator), and multi-agent coordination (Rainbow Teaming samvelyan2024rainbow, X-Teaming rahman2025x, EvoSynth chen2025evolve).
@@ 4.1 Experimental Setup
Our experimental setup is designed to ensure a rigorous and fair comparison against current state-of-the-art methods. To this end, we closely follow the evaluation protocols established by leading baseline frameworks, particularly X-Teaming (rahman2025x) and ActorAttack (ren2024derail). Following these works, we also use Harmbench Standardmazeika2024harmbench as the evaluation dataset. This dataset is designed to be comprehensive, with instructions balanced across 6 different risk categories specified in emerging AI regulation, including 6 semantic categories of behavior: Cybercrime & Unauthorized Intrusion, Chemical & Biological Weapons/Drugs, Misinformation & Disinformation, Harassment & Bullying, Illegal Activities, and General Harm, ensuring our evaluation represents a representative spectrum of potential harms.
@@ 4.1.1 Datasets and Models
For our experiments, we employ the HarmfulBench mazeika2024harmbench dataset, which comprises a curated collection of harmful queries designed to probe the safety vulnerabilities of MLLMs. We evaluate the performance of over 20 distinct target models, including MLLMs such as GPT-5.2 gpt5-2, GPT-5.1 gpt5, Claude Haiku 4.5 ClaudeHaiku45, Gemini 3 Pro Preview Gemini3ProPreview, Gemini 2.5 Flash Thinking comanici2025gemini, Mistral Large 3 Mistral32025, Llama-4-Maverick Meta2025Llama4, Llama-4-Scout Meta2025Llama4, Grok 4.1 Fast Grok41Fast2025, and Doubao-Seed-1.6 Seed16, as well as LLMs including Qwen3-Max qwen3max, Qwen3-235B-A22B-Thinking Qwen3Thinking235B, Qwen3-Next-80B-A3B Qwen2025Qwen3Next80B, DeepSeek-R1 guo2025deepseek, DeepSeek-V3.2 liu2025deepseek, Kimi K2-Instruct-0905 kimiteam2025kimik2openagentic, MiniMax-M2 MiniMax2025M2, GLM-4.6 vteam2025glm45vglm41vthinkingversatilemultimodal, Hunyuan-A13B-Instruct HunyuanA13BInstruct, and ERNIE-4.5-300B-A47B ernie2025technicalreport. These models represent the current frontier in AI safety and alignment, making them challenging targets.
@@ 4.1.3 Implementation Details
All experiments are conducted within a unified and reproducible OpenRT framework, utilizing a modular orchestration design to ensure fair comparisons across various attack methods and target models. We assign specialized models for distinct roles: all helper models (e.g., attacker, mutator, planner, optimizer) utilize DeepSeek-V3.2 liu2025deepseek with a temperature of 1.0 to encourage diverse and creative adversarial prompt generation, while the judge model employs GPT-4o-mini hurst2024gpt with a temperature of 0.0 to ensure deterministic and consistent safety evaluation, using a success threshold score of 5 for binary jailbreak classification. For attacks requiring semantic similarity computation, we leverage text-embedding-3-large as our embedding backbone. For multimodal attacks targeting vision-language models, we incorporate Imagen-4.0-fast as the diffusion-based image generator for synthesizing adversarial visual content, and Qwen2.5-VL-32B-Instruct serves as the vision-language attacker model for attacks requiring visual understanding and multimodal reasoning capabilities. Experiments are executed using a concurrent orchestration pipeline with 25 parallel workers per target model to maximize throughput while respecting API rate limits, and the evaluation module employs 32 parallel workers for efficient batch assessment of attack outcomes. All attack results are systematically logged with comprehensive metadata, including the full attack trajectory, intermediate prompts, model responses, and execution time for reproducibility analysis.

### v2

@@ 3.2.3 Attack
Table 2 classifies attacks by strategy paradigm, including gradient optimization, genetic algorithms, LLM-driven refinement, linguistic encoding, contextual deception, multimodal exploitation, and multi-agent coordination. White-box attacks include gradient-based methods like GCG Zou et al. (2023) and Visual Jailbreak Qi et al. (2024). Black-box attacks span multiple categories: optimization and fuzzing approaches (AutoDAN Liu et al. (2024a), GPTFuzzer Yu et al. (2023)), LLM-driven refinement (PAIR Chao et al. (2025), DrAttack Li et al. (2024c)), linguistic and encoding manipulations (CipherChat Yuan et al. (2024), CodeAttack Ren et al. (2024a), FlipAttack Liu et al. (2024b)), contextual deception strategies (Crescendo Russinovich et al. (2025), RedQueen Jiang et al. (2024b)), multimodal-specific attacks (FigStep Gong et al. (2025b), IDEATOR Wang et al. (2025a)), and multi-agent coordination (Rainbow Teaming Samvelyan et al. (2024), X-Teaming Rahman et al. (2025), EvoSynth Chen et al. (2025b)).
@@ 4.1 Experimental Setup
Our experimental setup is designed to ensure a rigorous and fair comparison against current state-of-the-art methods. To this end, we closely follow the evaluation protocols established by leading baseline frameworks, particularly X-Teaming (Rahman et al., 2025) and ActorAttack (Ren et al., 2024b). Following these works, we also use Harmbench StandardMazeika et al. (2024) as the evaluation dataset. This dataset is designed to be comprehensive, with instructions balanced across 6 different risk categories specified in emerging AI regulation, including 6 semantic categories of behavior: Cybercrime & Unauthorized Intrusion, Chemical & Biological Weapons/Drugs, Misinformation & Disinformation, Harassment & Bullying, Illegal Activities, and General Harm, ensuring our evaluation represents a representative spectrum of potential harms.
@@ 4.1.1 Datasets and Models
For our experiments, we employ the HarmfulBench Mazeika et al. (2024) dataset, which comprises a curated collection of harmful queries designed to probe the safety vulnerabilities of MLLMs. We evaluate the performance of over 20 distinct target models, including MLLMs such as GPT-5.2 OpenAI (2025b), GPT-5.1 OpenAI (2025a), Claude Haiku 4.5 Anthropic (2025), Gemini 3 Pro Preview Google (2025), Gemini 2.5 Flash Thinking Comanici et al. (2025), Mistral Large 3 AI (2025b), Llama-4-Maverick AI (2025a), Llama-4-Scout AI (2025a), Grok 4.1 Fast xAI (2025), and Doubao-Seed-1.6 Seed (2025), as well as LLMs including Qwen3-Max Team (2025c), Qwen3-235B-A22B-Thinking Team (2025b), Qwen3-Next-80B-A3B Team (2025d), DeepSeek-R1 Guo et al. (2025), DeepSeek-V3.2 Liu et al. (2025a), Kimi K2-Instruct-0905 Team (2025a), MiniMax-M2 MiniMax (2025), GLM-4.6 Team (2025e), Hunyuan-A13B-Instruct Tencent Hunyuan Team (2025), and ERNIE-4.5-300B-A47B Baidu-ERNIE-Team (2025). These models represent the current frontier in AI safety and alignment, making them challenging targets.
@@ 4.1.3 Implementation Details
All experiments are conducted within a unified and reproducible OpenRT framework, utilizing a modular orchestration design to ensure fair comparisons across various attack methods and target models. We assign specialized models for distinct roles: all helper models (e.g., attacker, mutator, planner, optimizer) utilize DeepSeek-V3.2 Liu et al. (2025a) with a temperature of 1.0 to encourage diverse and creative adversarial prompt generation, while the judge model employs GPT-4o-mini Hurst et al. (2024) with a temperature of 0.0 to ensure deterministic and consistent safety evaluation, using a success threshold score of 5 for binary jailbreak classification. For attacks requiring semantic similarity computation, we leverage text-embedding-3-large as our embedding backbone. For multimodal attacks targeting vision-language models, we incorporate Imagen-4.0-fast as the diffusion-based image generator for synthesizing adversarial visual content, and Qwen2.5-VL-32B-Instruct serves as the vision-language attacker model for attacks requiring visual understanding and multimodal reasoning capabilities. Experiments are executed using a concurrent orchestration pipeline with 25 parallel workers per target model to maximize throughput while respecting API rate limits, and the evaluation module employs 32 parallel workers for efficient batch assessment of attack outcomes. All attack results are systematically logged with comprehensive metadata, including the full attack trajectory, intermediate prompts, model responses, and execution time for reproducibility analysis.